the product of two consecutive positive even integers is 288. Find the integers.
Answer by malakumar_kos@yahoo.com(315)


the product of two consecutive positive even integers is 288. Find the integers
Let the first even number be = x
the second even number is = x+2
product of even integers = 288
x(x+2) = 288
x^2+2x-288 = 0
x^2+18x-16x-288 = 0
x(x+18)-16(x+18) = 0
(x+18).(x-16) = 0
since the numbers are positive integers we
consider only (x-16) = 0
x = 16
x+2 = 16+2 = 18
The two numbers (even consecutive) are 16 and 18.